# Semi-Random Graphs, Robust Asymmetry, and Reconstruction

### v1

@@ Abstract
In particular, Bollobás (1990) demonstrated that almost all Erdős-Rényi random graphs $G=(V,E)\sim\mathscr{G}(n,p)$ enjoy the property that their induced subgraphs on $n-\Theta(1)$ vertices are asymmetric and mutually non-isomorphic, for $1-p,p=\Omega(\log(n)/n)$. As our primary result, we demonstrate that this property is robust against perturbation—even when an adversary is permitted to add/remove each vertex pair in $V^{(2)}$ with (independent) arbitrarily large constant probability. Exploiting this result, we derive asymptotic characterizations of asymmetry in random graphs with large planted structure and bounded adversarial corruptions, along with improved bounds on the probability mass of nonreconstructible graphs in $\mathscr{G}(n,p)$.
@@ 1 Introduction
The Graph Reconstruction Conjecture, or simply the Reconstruction Conjecture, dates to the work of Ulam (1960) and Kelly (1942), and remains one of the foremost open problems in graph theory. Informally, the conjecture states that each undirected graph on at least three vertices is reconstructible, meaning it is determined up to isomorphism by its multiset of unlabeled induced subgraphs.
@@
Note that Question ($i.$) commences with the worst-case perspective of the Reconstruction Conjecture by fixing an underlying graph $G$ — thought of as adversarially selected — yet mollifies some of its structure by way of random edge noise. Patently, this perspective is motivated by the smoothed analysis of Spielman and Teng (2004), which measures the complexity of algorithms on worst-case inputs that experience slight random perturbations before their runtimes are measured. In the context of graph reconstruction, this outlook can be thought of as measuring the brittleness or sparsity of nonreconstructible graphs; in order for an adversary to succeed, they must select a graph whose nonreconstructibility is unlikely to be ameliorated by edge noise.
By a similar token, Question (ii.ii.) modifies the standard probabilistic treatment of graph reconstruction by endowing it with a component of worst-case analysis. Namely, an adversary is granted the opportunity to adjust each random graph before examining its reconstructibility, potentially allowing them to destroy brittle graph properties previously used to demonstrate reconstrucibility (e.g., subgraph asymmetry). This perspective is inspired by the analysis of algorithms in the semi-random setting (Blum and Spencer, 1995; Feige and Kilian, 2001), in which inputs are drawn according to a natural distribution and subsequently modified by an adversary with constrained editing power.
@@ 1.1 Primary Results and Technical Overview
Intuitively, should we allow an adversary to edit a limited number of edges anywhere in $G$, we hope that all graphs in $B_{\varepsilon}(G)$ be reconstructible. Should we constrain an adversary’s changes to a (possibly random) set of pairs $S$, we hope that all graphs in $B_{S}(G)$ be reconstructible. To start, we find it most fruitful to uniformly draw a subset $S\subset V^{(2)}$ of order $\varepsilon=\varepsilon(n)$ and examine the behavior of $B_{S}(G)$. Note that for appropriate $\varepsilon\in\Omega(n^{2}\log(n))$, an asymptotic reconstructibility result for $B_{S}(G)$ would suffice to prove the reconstruction conjecture for large $n$. This observation furnishes another measure of progress towards the full conjecture: the largest $\varepsilon(n)$ such that all graphs in $B_{S}(G)$ remain reconstructible asymptotically. Loosely, this formalization of the semi-random model can be seen as, instead of uniformly widening the hypercube bound on $\textbf{{p}}\in[0,1]^{N}$ (as is our objective under the smoothed interpretation), allowing a random (or adversarially chosen) subset of the entries in this tuple to be arbitrary. Studying these settings, we give strengthened and refined forms of existing probabilistic graph reconstruction results. We summarize our contributions as follows.
@@
Let $\delta\in\mathbb{N}$, $c\in(0,1)$, $\beta:\mathbb{N}\to(0,1)$, and $\varepsilon:\mathbb{N}\to\mathbb{N}$ such that $1-\beta,\beta\in\Omega(\log(n)/n)$ and $\varepsilon=\varepsilon(n)\leq cn^{2}$. There exists $\alpha>0$ such that if $\textbf{{p}}(n)\in[\alpha\beta(n),1-\alpha\beta(n)]^{\binom{n}{2}}$, then for $G\sim\mathscr{G}(n,\textbf{{p}})$ and independent uniformly drawn subset $S\subset V^{(2)}$ of order $\varepsilon$, it holds with failure probability at most $\exp(-\Omega(\beta(n)n))$ that for all $H\in B_{S}(G)$, all induced subgraphs of $H$ on $n-\delta$ vertices are asymmetric and mutually non-isomorphic.
We remark that for $\beta(n)\in o(1)$ and fixed $\rho>0$, we can take $\alpha>0$ so that the failure probability is bounded by $\exp(-\rho\beta(n)n)$. Up to constants in the bound on $p(n)$, the asymmetry result of Bollobás (1990) can be seen as the special case where $\varepsilon=0$.
@@
When $G\sim\mathscr{G}(n,1/2)$ and no edits are allowed, enumerative techniques similar to those of Müller (1977) give a failure probability in Theorem 1.1 of $\exp(-\Omega(n))$. If we want to ensure asymmetry on all graphs $G^{\prime}\in B_{\varepsilon}(G)$, we can simply observe an $O(n^{2\varepsilon})$ cardinality blowup, so it is enough to take $\varepsilon\leq Bn/\log(n)$ for carefully chosen $B>0$. This does not translate nicely to the non-uniform setting, however (e.g. $p\neq 1/2$). Using Theorem 1.1, we obtain a strict generalization of this observation via conditioning. We obtain even stronger results when adversarial perturbations are restricted to (or planted on) fixed subsets of $V^{(2)}$.
@@
For $|\textbf{{p}}|_{\infty}\in\Theta(1)$, our result is tight up to constants; trivially, one can plant 2n32n-3 nonedges to guarantee the existence of a chosen transposition on any sampled graph. As a concrete illustration, when $p\in\Theta(1)$, one preserves asymmetry even when a clique is planted on $\Theta(\sqrt{n})$ vertices. In contrast, the clique number of $\mathscr{G}(n,p)$ is known to be $\Theta(\log n)$ asymptotically almost surely (Grimmett and McDiarmid, 1975).
@@
One can view Corollary 1.3 as establishing robustness against perturbations made after the graph has been sampled. Though Corollary 1.3 admits logarithmically smaller $\varepsilon$ than the results of Erdős and Rényi (1963); Kim et al. (2002), and does not provide tight characterizations of resilience, we address the nonexistence of isomorphisms between proper subgraphs and allow for non-uniformity in edge probabilities, which rules out exact computation. [^5] Note that Corollary 1.3 devolves into a global asymmetry result when $\delta=1$; note that all subgraphs on n1n-1 vertices being mutually non-isomorphic is sufficient for global asymmetry. This latter point enables the smoothed setting, in which edge probabilities may differ by 1o(1)1-o(1).
@@
While our asymmetry results may be of independent interest, the consequences in our setting are immediate. Taking $\delta=2$, the conclusions of Theorems 1.1, 1.2 and 1.3 are sufficient for reconstructibility via Lemma 3.1. This yields reconstruction results for random graphs with: 1) adversarial edge edits constrained to a random subset of $O(n^{2})$ vertex pairs, 2) planted structure on $O(\beta(n)n)$ fixed vertex pairs, or 3) at most $O(\beta(n)n/\log(n))$ adversarial edge edits anywhere in the graph. We state these results in full generality in Section 3.3. After providing some notation, we state special cases here.
@@
When each edge is drawn with uniform probability $p\in\Theta(\log(n)/n)$, the result of Bollobás (1990) guarantees convergence to zero only for $\varepsilon=0$. Corollary 1.5 demonstrates that in this same regime, we can in fact take any constant $\varepsilon$. Using techniques from Bollobás (1990), one can show that the same results hold if we instead let $\mathfrak{R}_{n}$ denote the family of graphs on $n$ vertices with reconstruction number three. [^6] A graph is said to have reconstruction number $k\leq n$ if it determined by some order-$k$ collection of its induced subgraphs (Harary and Plantholt, 1985).
@@ 1.2 Related Work
In the present work, we are concerned with the Vertex Reconstruction Conjecture. Graph families proven to be reconstructible in this setting include regular graphs, disconnected graphs, and trees (Kelly, 1957). More recently, Farhadian (2017) observed connections between the reconstructibility of a graph and the uniqueness and asymmetry of its large subgraphs. On the empirical front, McKay (1997, 2022) verified that all graphs up to 13 vertices are reconstructible. For an overview of standard results and alternative variants of the conjecture, see Bondy (1991).
@@
The probabilistic examination of graph reconstruction dates to Müller (1976), who proved that for any $\alpha\in(0,1)$, almost all graphs are reconstructible from their multiset of subgraphs on $\frac{n}{2}(1+\alpha)$ vertices. Using an enumerative argument, Müller proved that for $G\sim\mathscr{G}(n,1/2)$, it holds with probability 1o(1)1-o(1) that all sufficiently large subgraphs of $G$ must be asymmetric and mutually non-isomorphic, a sufficient condition for reconstructibility. The seminal work of Bollobás (1990) refined the analysis for $\mathscr{G}(n,p)$ with $p\in\Omega(\log(n)/n)$, proving that asymptotically almost surely $G\sim\mathscr{G}(n,p)$ is reconstructible from any three subgraphs on n1n-1 vertices. More recently, Spinoza and West (2019) used the asymmetry result of Müller (1976) to prove that for $\ell\leq\frac{n}{2}(1-\alpha)$, almost all graphs are reconstructible from some multiset of $\binom{\ell+2}{2}$ subgraphs on $n$-$\ell$ vertices.
@@
The literature on graph asymmetry is even more extensive. Erdős and Rényi (1963) initiated the study of this notion by proving that almost all graphs are asymmetric. In particular, they proved that asymptotically almost surely, one must remove and add a total of at least $\frac{n}{2}(1-\alpha)$ edges to make $G\sim\mathscr{G}(n,1/2)$ symmetric. In recent years, this result has joined the literature on property resilience (Sudakov and Vu, 2008; Frieze and Karoński, 2015; Bollobás, 2011). More generally, it is folklore that for $1-p,p\geq\log(n)/n$, one must asymptotically almost surely add and remove at least (2+o(1))np(1p)(2+o(1))np(1-p) edges somewhere in $G\sim\mathscr{G}(n,p)$ to achieve symmetry. Kim et al. (2002) refined this result by proving that for the same range of $p$, one must asymptotically almost surely add and remove at least (1+o(1))np(1p)(1+o(1))np(1-p) edges adjacent to a single vertex. A number of works have also studied asymmetry in different graph distributions (Kim et al., 2002; Bollobás, 1982; McKay and Wormald, 1984; O’Donnell et al., 2014). The work of O’Donnell et al. (2014) comes closest to ours in terminology; examining hardness in a robust variant of the Graph Isomorphism problem, they prove that the uniform distribution over graphs on $n$ vertices and $m\in\Omega(n)$ edges satisfies a strengthened notion of global asymmetry that decays with $n$ if $m$ is super-linear.
@@
Also related to the sensitivity of graph asymmetry is the work of Brewer et al. (2018) and Hartke et al. (2009). These works examine changes in the graph automorphism group under addition and deletion of carefully chosen edges. Brewer et al. (2018) take on a dual perspective to that of Erdős and Rényi (1963), studying the minimum number of edges one must add and delete in order to make a graph asymmetric. Hartke et al. (2009) prove that for any two finite groups, there exists a graph from which one can remove a vertex so that the automorphism group passes between these groups. This result can be seen as a barrier to proving the Reconstruction Conjecture via characterization of subgraph symmetries.
@@ 2.2 Graph Reconstruction
We direct our attention exclusively to finite undirected graphs; for an introduction to standard notation, see Bollobás (2002, 2011). For a graph $G=(V,E)$, we denote $v(G)=|V|$ and $e(G)=|E|$. For a set of vertices $U\subseteq V$, $G[U]$ denotes the induced subgraph of $G$ on $U$, i.e., $G[U]=(U,\{e\in E:e\subset U\})$. For a vertex $v\in V$, we let $G-v=G[V\setminus\{v\}]$. More generally, for any subset $U\subset V$, we let $G-U=G[V\setminus U]$. Analogously, for any subset $F\subset V^{(2)}$, we let $G+F=(V,E\cup F)$ and $G-F=(V,E\setminus F)$. Henceforth, for concision, we refer to induced subgraphs simply as subgraphs.
A graph isomorphism $\varphi\colon G\xrightarrow{\cong}H$ is a bijection $\varphi\colon V(G)\to V(H)$ such that $xy\in E(G)$ if and only if $\varphi(x)\varphi(y)\in E(H)$. If there exists an isomorphism $\varphi\colon G\xrightarrow{\cong}H$, we say that $G$ and $H$ are isomorphic and write $G\cong H$. A graph automorphism on $G$ is an isomorphism $\varphi\colon G\xrightarrow{\cong}G$. We let $\operatorname{Aut}G$ denote the set of all automorphism on $G$. The identity is always an automorphism; if no other automorphism exists, we call $G$ asymmetric. If $G$ is not asymmetric, we call it symmetric. We say that two vertices $x,y\in V(G)$ are similar, and write $x\sim y$, if there exists $\varphi\in\operatorname{Aut}G$ such that $\varphi(x)=y$. We say that a set of vertices $U\subset V(G)$ is fixed in $G$ if $\varphi(U)=U$ for all $\varphi\in\operatorname{Aut}G$. We say a subgraph of $G$ is unique if it is isomorphic to no other subgraph of $G$.
@@
Conjecture 2.1 (Ulam (1960); Kelly (1942)).
@@
In general, we say that a graph $G$ is reconstructible if whenever $G\sim H$, it holds that $G\cong H$. A family of graphs $\mathcal{C}$ is reconstructible if all $G\in\mathcal{C}$ are reconstructible. We let $\mathfrak{R}_{n}$ denote the family of reconstructible graphs on $n$ vertices, and $\overline{\mathfrak{R}}_{n}$ the family of nonreconstructible graphs on $n$ vertices.

### v2

@@ Abstract
In particular, bollobas1990almost demonstrated that almost all Erdős-Rényi random graphs $G=(V,E)\sim\mathscr{G}(n,p)$ enjoy the property that their induced subgraphs on $n-\Theta(1)$ vertices are asymmetric and mutually non-isomorphic, for $1-p,p=\Omega(\log(n)/n)$. As our primary result, we demonstrate that this property is robust against perturbation—even when an adversary is permitted to add/remove each vertex pair in $V^{(2)}$ with (independent) arbitrarily large constant probability. Exploiting this result, we derive asymptotic characterizations of asymmetry in random graphs with large planted structure and bounded adversarial corruptions, along with improved bounds on the probability mass of nonreconstructible graphs in $\mathscr{G}(n,p)$.
@@ 1 Introduction
The Graph Reconstruction Conjecture, or simply the Reconstruction Conjecture, dates to the work of ulam1960collection and kelly1942isometric, and remains one of the foremost open problems in graph theory. Informally, the conjecture states that each undirected graph on at least three vertices is reconstructible, meaning it is determined up to isomorphism by its multiset of unlabeled induced subgraphs.
@@
Note that Question ($i.$) commences with the worst-case perspective of the Reconstruction Conjecture by fixing an underlying graph $G$ — thought of as adversarially selected — yet mollifies some of its structure by way of random edge noise. Patently, this perspective is motivated by the smoothed analysis of spielman2004smoothed, which measures the complexity of algorithms on worst-case inputs that experience slight random perturbations before their runtimes are measured. In the context of graph reconstruction, this outlook can be thought of as measuring the brittleness or sparsity of nonreconstructible graphs; in order for an adversary to succeed, they must select a graph whose nonreconstructibility is unlikely to be ameliorated by edge noise.
By a similar token, Question (ii.ii.) modifies the standard probabilistic treatment of graph reconstruction by endowing it with a component of worst-case analysis. Namely, an adversary is granted the opportunity to adjust each random graph before examining its reconstructibility, potentially allowing them to destroy brittle graph properties previously used to demonstrate reconstrucibility (e.g., subgraph asymmetry). This perspective is inspired by the analysis of algorithms in the semi-random setting (blum2995coloring; feige2001heuristics), in which inputs are drawn according to a natural distribution and subsequently modified by an adversary with constrained editing power.
@@ 1.1 Primary Results and Technical Overview
Intuitively, should we allow an adversary to edit a limited number of edges anywhere in $G$, we hope that all graphs in $B_{\varepsilon}(G)$ be reconstructible. Should we constrain an adversary’s changes to a (possibly random) set of pairs $S$, we hope that all graphs in $B_{S}(G)$ be reconstructible. To start, we find it most fruitful to uniformly draw a subset $S\subset V^{(2)}$ of order $\varepsilon=\varepsilon(n)$ and examine the behavior of $B_{S}(G)$. Note that for appropriate $\varepsilon\in\Omega(n^{2}\log(n))$, an asymptotic reconstructibility result for $B_{S}(G)$ would suffice to prove the Reconstruction Conjecture for sufficiently large $n$ (see Appendix C). This observation furnishes another measure of progress towards the full conjecture: the largest $\varepsilon(n)$ such that all graphs in $B_{S}(G)$ remain reconstructible asymptotically. Loosely, this formalization of the semi-random model can be seen as, instead of uniformly widening the hypercube bound on $\textbf{{p}}\in[0,1]^{N}$ (as is our objective under the smoothed interpretation), allowing a random (or adversarially chosen) subset of the entries in this tuple to be arbitrary. Studying these settings, we give strengthened and refined forms of existing probabilistic graph reconstruction results. We summarize our contributions as follows.
@@
Let $\delta\in\mathbb{N}$, $c\in(0,1/2)$, $\beta:\mathbb{N}\to(0,1)$, and $\varepsilon:\mathbb{N}\to\mathbb{N}$ such that $1-\beta,\beta\in\Omega(\log(n)/n)$ and $\varepsilon=\varepsilon(n)\leq cn^{2}$. There exists $\alpha>0$ such that if $\textbf{{p}}(n)\in[\alpha\beta(n),1-\alpha\beta(n)]^{\binom{n}{2}}$, then for $G\sim\mathscr{G}(n,\textbf{{p}})$ and independent uniformly drawn subset $S\subset V^{(2)}$ of order $\varepsilon$, it holds with failure probability at most $\exp(-\Omega(\beta(n)n))$ that for all $H\in B_{S}(G)$, all induced subgraphs of $H$ on $n-\delta$ vertices are asymmetric and mutually non-isomorphic.
We remark that for $\beta(n)\in o(1)$ and fixed $\rho>0$, we can take $\alpha>0$ so that the failure probability is bounded by $\exp(-\rho\beta(n)n)$. Up to constants in the bound on $p(n)$, the asymmetry result of bollobas1990almost can be seen as the special case where $\varepsilon=0$.
@@
When $G\sim\mathscr{G}(n,1/2)$ and no edits are allowed, enumerative techniques similar to those of muller1977edge give a failure probability in Theorem 1.1 of $\exp(-\Omega(n))$. If we want to ensure asymmetry on all graphs $G^{\prime}\in B_{\varepsilon}(G)$, we can simply observe an $O(n^{2\varepsilon})$ cardinality blowup, so it is enough to take $\varepsilon\leq Bn/\log(n)$ for carefully chosen $B>0$. This does not translate nicely to the non-uniform setting, however (e.g. $p\neq 1/2$). Using Theorem 1.1, we obtain a strict generalization of this observation via conditioning. We obtain even stronger results when adversarial perturbations are restricted to (or planted on) fixed subsets of $V^{(2)}$.
@@
For $|\textbf{{p}}|_{\infty}\in\Theta(1)$, our result is tight up to constants; trivially, one can plant 2n32n-3 nonedges to guarantee the existence of a chosen transposition on any sampled graph. As a concrete illustration, when $p\in\Theta(1)$, one preserves asymmetry even when a clique is planted on $\Theta(\sqrt{n})$ vertices. In contrast, the clique number of $\mathscr{G}(n,p)$ is known to be $\Theta(\log n)$ asymptotically almost surely (grimmett1975colouring).
@@
One can view Corollary 1.3 as establishing robustness against perturbations made after the graph has been sampled. Though Corollary 1.3 admits logarithmically smaller $\varepsilon$ than the results of erdos1963asymmetric; kim2002asymmetry, and does not provide tight characterizations of resilience, we address the nonexistence of isomorphisms between proper subgraphs and allow for non-uniformity in edge probabilities, which rules out exact computation. [^5] Note that Corollary 1.3 devolves into a global asymmetry result when $\delta=1$; note that all subgraphs on n1n-1 vertices being mutually non-isomorphic is sufficient for global asymmetry. This latter point enables the smoothed setting, in which edge probabilities may differ by 1o(1)1-o(1).
@@
While our asymmetry results may be of independent interest, the consequences in our setting are immediate. Taking $\delta=2$, the conclusions of Theorems 1.1, 1.2 and 1.3 are sufficient for reconstructibility via Corollary 3.2. This yields reconstruction results for random graphs with: 1) adversarial edge edits constrained to a random subset of $O(n^{2})$ vertex pairs, 2) planted structure on $O(\beta(n)n)$ fixed vertex pairs, or 3) at most $O(\beta(n)n/\log(n))$ adversarial edge edits anywhere in the graph. We state these results in full generality in Section 3.3. After providing some notation, we state special cases here.
@@
When each edge is drawn with uniform probability $p\in\Theta(\log(n)/n)$, the result of bollobas1990almost guarantees convergence to zero only for $\varepsilon=0$. Corollary 1.5 demonstrates that in this same regime, we can in fact take any constant $\varepsilon$. Using techniques from bollobas1990almost, one can show that the same results hold if we instead let $\mathfrak{R}_{n}$ denote the family of graphs on $n$ vertices with reconstruction number three. [^6] A graph is said to have reconstruction number $k\leq n$ if it determined by some order-$k$ collection of its induced subgraphs (harary1985reconstruction).
@@ 1.2 Related Work
In the present work, we are concerned with the Vertex Reconstruction Conjecture. Graph families proven to be reconstructible in this setting include regular graphs, disconnected graphs, and trees (kelly1957congruence). More recently, farhadian2017simple observed connections between the reconstructibility of a graph and the uniqueness and asymmetry of its large subgraphs. On the empirical front, mckay1997small; mckay2022reconstruction verified that all graphs up to 13 vertices are reconstructible. For an overview of standard results and alternative variants of the conjecture, see bondy1991manual.
@@
The probabilistic examination of graph reconstruction dates to muller1976probabilistic, who proved that for any $\alpha\in(0,1)$, almost all graphs are reconstructible from their multiset of subgraphs on $\frac{n}{2}(1+\alpha)$ vertices. Using an enumerative argument, Müller proved that for $G\sim\mathscr{G}(n,1/2)$, it holds with probability 1o(1)1-o(1) that all sufficiently large subgraphs of $G$ must be asymmetric and mutually non-isomorphic, a sufficient condition for reconstructibility. The seminal work of bollobas1990almost refined the analysis for $\mathscr{G}(n,p)$ with $p\in\Omega(\log(n)/n)$, proving that asymptotically almost surely $G\sim\mathscr{G}(n,p)$ is reconstructible from any three subgraphs on n1n-1 vertices. More recently, spinoza2019reconstruction used the asymmetry result of muller1976probabilistic to prove that for $\ell\leq\frac{n}{2}(1-\alpha)$, almost all graphs are reconstructible from some multiset of $\binom{\ell+2}{2}$ subgraphs on $n$-$\ell$ vertices.
@@
The literature on graph asymmetry is even more extensive. erdos1963asymmetric initiated the study of this notion by proving that almost all graphs are asymmetric. In particular, they proved that asymptotically almost surely, one must remove and add a total of at least $\frac{n}{2}(1-\alpha)$ edges to make $G\sim\mathscr{G}(n,1/2)$ symmetric. In recent years, this result has joined the literature on property resilience (sudakov2008local; frieze2015introduction; bollobas2011random). More generally, it is folklore that for $1-p,p\geq\log(n)/n$, one must asymptotically almost surely add and remove at least (2+o(1))np(1p)(2+o(1))np(1-p) edges somewhere in $G\sim\mathscr{G}(n,p)$ to achieve symmetry. kim2002asymmetry refined this result by proving that for the same range of $p$, one must asymptotically almost surely add and remove at least (1+o(1))np(1p)(1+o(1))np(1-p) edges adjacent to a single vertex. A number of works have also studied asymmetry in different graph distributions (kim2002asymmetry; bollobas1982asymptotic; mckay1984automorphisms; odonnell2014hardness). The work of odonnell2014hardness comes closest to ours in terminology; examining hardness in a robust variant of the Graph Isomorphism problem, they prove that the uniform distribution over graphs on $n$ vertices and $m\in\Omega(n)$ edges satisfies a strengthened notion of global asymmetry that decays with $n$ if $m$ is super-linear.
@@
Also related to the sensitivity of graph asymmetry is the work of brewer2018asymmetric and hartke2009automorphism. These works examine changes in the graph automorphism group under addition and deletion of carefully chosen edges. brewer2018asymmetric take on a dual perspective to that of erdos1963asymmetric, studying the minimum number of edges one must add and delete in order to make a graph asymmetric. hartke2009automorphism prove that for any two finite groups, there exists a graph from which one can remove a vertex so that the automorphism group passes between these groups. This result can be seen as a barrier to proving the Reconstruction Conjecture via characterization of subgraph symmetries.
@@ 2.2 Graph Reconstruction
We direct our attention exclusively to finite undirected graphs; for an introduction to standard notation, see bollobas2002modern; bollobas2011random. For a graph $G=(V,E)$, we denote $v(G)=|V|$ and $e(G)=|E|$. For a set of vertices $U\subseteq V$, $G[U]$ denotes the induced subgraph of $G$ on $U$, i.e., $G[U]=(U,\{e\in E:e\subset U\})$. For a vertex $v\in V$, we let $G-v=G[V\setminus\{v\}]$. More generally, for any subset $U\subset V$, we let $G-U=G[V\setminus U]$. Analogously, for any subset $F\subset V^{(2)}$, we let $G+F=(V,E\cup F)$ and $G-F=(V,E\setminus F)$. Henceforth, for concision, we refer to induced subgraphs simply as subgraphs.
A graph isomorphism $\varphi\colon G\xrightarrow{\cong}H$ is a bijection $\varphi\colon V(G)\to V(H)$ such that $xy\in E(G)$ if and only if $\varphi(x)\varphi(y)\in E(H)$. If there exists an isomorphism $\varphi\colon G\xrightarrow{\cong}H$, we say that $G$ and $H$ are isomorphic and write $G\cong H$. A graph automorphism on $G$ is an isomorphism $\varphi\colon G\xrightarrow{\cong}G$. We let $\operatorname{Aut}G$ denote the set of automorphisms on $G$. The identity is always an automorphism; if no other automorphism exists, we call $G$ asymmetric. If $G$ is not asymmetric, we call it symmetric. We say that two vertices $x,y\in V(G)$ are similar, and write $x\sim y$, if there exists $\varphi\in\operatorname{Aut}G$ such that $\varphi(x)=y$. We say that a set of vertices $U\subset V(G)$ is fixed in $G$ if $\varphi(U)=U$ for all $\varphi\in\operatorname{Aut}G$. We say a subgraph of $G$ is unique if it is isomorphic to no other subgraph of $G$.
@@
Conjecture 2.1 (ulam1960collection; kelly1942isometric).
@@
In general, we say that a graph $G$ is reconstructible if whenever $G\sim H$, it holds that $G\cong H$. A family of graphs $\mathcal{C}$ is reconstructible if all $G\in\mathcal{C}$ are reconstructible. We let $\mathfrak{R}_{n}$ denote the family of reconstructible graphs on $n$ vertices, and $\overline{\mathfrak{R}}_{n}$ the family of nonreconstructible graphs on $n$ vertices. A classical tool in graph reconstruction is Kelly’s lemma, which we state below alongside two relevant facts.
Lemma 2.2 (kelly1957congruence).
Suppose $G\sim H$. Then for any graph $F$ on at most n1n-1 vertices, $G$ and $H$ have the same number of subgraphs isomorphic to $F$.
Fact 2.3.
Let $\sigma\colon G\xrightarrow{\sim}H$, and $U\subset V(G)$ such that $G[U]$ is a unique subgraph of $G$. Then $G[U]\cong H[\sigma(U)]$, and $H[\sigma(U)]$ is unique in $H$.
Proof.
Let $\varphi_{x}:G-x\xrightarrow{\cong}H-\sigma(x)$ denote isomorphisms given by the existence of $\sigma$. By Kelly’s lemma, there is a subset $U^{\prime}\subset V(H)$ such that $H[U^{\prime}]\cong G[U]$ and $H[U^{\prime}]$ is unique in $H$. Also note that for all $x\in V\setminus U$, $\varphi_{x}$ restricts to an isomorphism $G[U]\xrightarrow{\cong}H[\varphi_{v}(U)]$, so $H[\varphi_{v}(U)]$ is a subgraph of $H$ isomorphic to $G[U]$, and $\varphi_{v}(U)=U^{\prime}$. In particular, $\sigma(x)\notin U^{\prime}$ for all $x\in V\setminus U$, and $\sigma(U)=U^{\prime}$. ∎